OpenStudy (anonymous):

The three sides of an equilateral triangle measure x + 12, 2x - 5 and 3x - 22 inches. The sides of the triangle measure 17 inches 29 inches 32.5 inches 44.5 inches

Parth (parthkohli):

Equilateral triangle means that all sides are equal. We can do the same thing in this. Just put up an equation.

Parth (parthkohli):

Take any two sides and then equate them

OpenStudy (turingtest):

@LovelyMegan did you understand Parth's explanation?

OpenStudy (anonymous):

Explain more plz

OpenStudy (anonymous):

is it the same formula?

Parth (parthkohli):

x + 12 = 2x - 5 Solve this

Parth (parthkohli):

Alternatively, you can do this too, since they're all equal. 2x - 5 = 3x - 22

Parth (parthkohli):

Or even this. 3x - 22 = x + 12

OpenStudy (turingtest):

@LovelyMegan have you tried any of these formulas yet?

OpenStudy (anonymous):

just about to i had to take a quick break

OpenStudy (turingtest):

well please let me know if/where you get stuck in solving this problem :)

OpenStudy (anonymous):

which formula am i use n ?

OpenStudy (anonymous):

either?

OpenStudy (turingtest):

any one of the three Parth gave you; they all give the same answer. Why? Because the triangle is \(equilateral\) which means all sides have the same length, so we can set any two equal to solve for x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so im solve n for x?

OpenStudy (turingtest):

yep

OpenStudy (turingtest):

then you need to plug that value of x into any one of the expressions for the length of the side to get the actual answer

OpenStudy (anonymous):

wait but what is x

OpenStudy (turingtest):

that's your job: to find x with any of the three formulas parth gave you

Parth (parthkohli):

x is just a variable. When you find x, then plug it in.

Parth (parthkohli):

Make sure that when you plug-in, you get all of them the same.

OpenStudy (anonymous):

Uhhh im not getting any of these answer

OpenStudy (turingtest):

ok, which equation did you use?

OpenStudy (anonymous):

3x-22=x+12

OpenStudy (turingtest):

can you show me how you worked it? what you got for x ?

OpenStudy (anonymous):

 see /7 now im suck